A survey asked people how many phone numbers they have stored in their cell phone. The results are shown in the table.
Phone numbers Frequency
0-5 25
6-10 50
11-15 48
16-20 22
21-25 14
What is the probability that a person surveyed, chosen at random, has 6-10 phone numbers stored in his or her cell phone?

Answer by stanbon(75887)

A survey asked people how many phone numbers they have stored in their cell phone. The results are shown in the table.
Phone numbers Frequency
0-5 25
6-10 50
11-15 48
16-20 22
21-25 14
What is the probability that a person surveyed, chosen at random, has 6-10 phone numbers stored in his or her cell phone?
------
Ans:: 50/(25+50+48+22+14) = 50/155 = 10/31 = 0.3226
--------------
Cheers,
Stan H.
------------